OpenStudy (anonymous):

Write the equation of the line that is parallel to the line 3x - y = -3 and passes through the point (4, -2). y = -one thirdx - 6 y = -one thirdx - 14 y = 3x - 6 y = 3x - 14

OpenStudy (anonymous):

@bibby

OpenStudy (bibby):

get 3x - y = -3 into y = mx +b form. The slope of our line will be the same slope as that

OpenStudy (anonymous):

please give me a sec.

OpenStudy (anonymous):

y=3x+3

OpenStudy (bibby):

so what is the slope of the line we want?

OpenStudy (anonymous):

3

OpenStudy (bibby):

do now we're torn between c and d. I think we have to use the point slope formula to find out which

OpenStudy (bibby):

\[y-y_1=m(x-x_1)\] where y1 and x1 are the point (4,-2)

OpenStudy (anonymous):

y-(-2)=3(x-4)

OpenStudy (bibby):

y+2=3(x-4) solve for y

OpenStudy (anonymous):

y=3x−14

OpenStudy (bibby):

yeah, good job

OpenStudy (anonymous):

thanks do u mind helping me on 3 more questions if you can

OpenStudy (bibby):

I'll be away for 2ish minutes but tag me in your other questions

OpenStudy (anonymous):

ok thanks